# RealHiTBench: A Comprehensive Realistic Hierarchical Table Benchmark for Evaluating LLM-Based Table Analysis

### v1

@@ 1 Introduction
In recent years, there has been a lot of work on table analysis (Li et al., 2024b; Zhao et al., 2024). Specifically, an increasing number of studies discuss table question answering (TableQA) tasks. (Yang et al., 2024b; Sarkar and Lausen, 2023; Zhou et al., 2024; Nguyen et al., 2024; Li et al., 2024a). Additionally, the introduction of large language models (LLMs), including close-source models such as GPT-4o (OpenAI, 2023) and Gemini-1.5-Pro (Anil et al., 2023), and table-oriented models like TableGPT (Su et al., 2024) further enhances models’ capability of table comprehension. Given these advancements, the rapid development of LLMs and the continuous emergence of diverse tabular data in real-world scenarios have highlighted the need for a comprehensive evaluation of LLMs’ table understanding capabilities, putting forward new requirements on the relevant benchmarks.
@@
Unfortunately, we find that the current tabular data benchmarks, such as TAT-QA (Zhu et al., 2021), TableBench (Wu et al., 2024), and InfiAgent-DABench (Hu et al., 2024), largely consist of flat tables. More concretely, in such tables, each column represents an attribute, each row represents a record, and all data are stored in a simple one-dimensional format; see Figure 13 for visual demonstration. However, in practical applications, humans often organize relatively complex tables to represent multifaceted relationships between variables. Such hierarchically structured tables are popular in various domains, including economy, science, and employment, on public data platforms.
To address this issue, some benchmarks have considered hierarchical tables, such as HiTab (Cheng et al., 2022) and SciTab (Lu et al., 2023). However, these benchmarks fail to comprehensively present the LLMs’ understanding capabilities of complex table structures. For example, SciTab (Lu et al., 2023) and AIT-QA (Katsis et al., 2022) only consider tables of the scientific and airline domains, respectively. MM-Tab (Zheng et al., 2024) provides only image-based input, while hierarchical tables in realistic applications are also presented in a textual way. SpreadSheetBench (Ma et al., 2024) focuses on the operation of tables. HiTab (Cheng et al., 2022) organizes its tables with lossy JSON format, only contains fundamental QA tasks, and provides incomplete supervision. More importantly, most of these benchmarks focus primarily on relatively simple hierarchical tables, particularly those with a basic column hierarchy, where the hierarchy typically does not exceed two levels. To date, there is still a lack of dedicated benchmarks for comprehensively evaluating LLMs’ ability to understand complex tabular hierarchies.
@@ 2 Related Work
Table analysis is pivotal across numerous domains (He et al., 2024), while table question answering (TableQA) can effectively assess analytical capabilities (Müller et al., 2019). Recently, there has been a growing number of studies focusing on the ability of LLMs to comprehend tabular data (Singha et al., 2023; Nguyen et al., 2024; Deng et al., 2024). However, the development of LLMs leads to the emergence of different branches. Common tabular data are in text form to be processed by LLMs (Anil et al., 2023; Yang et al., 2024a; Dubey et al., 2024), while other image-based tables are suited to be processed by MLLMs (Wang et al., 2024; Liu et al., 2024). Additionally, some table-oriented LLMs for both modalities have also been proposed (Su et al., 2024; Zheng et al., 2024). Therefore, the flourishing development of LLMs makes challenging benchmarks for tabular data increasingly important.
@@
An increasing number of benchmarks for table analysis are discussed (Chen et al., 2020; Nan et al., 2022). However, a proportion of benchmarks focus more on other aspects, such as reasoning methods and supervision fine-tuning, while overlooking the structural complexity of the data, especially the table dimensions (Wu et al., 2024; Zheng et al., 2024). Although some recent benchmarks have introduced the concepts of hierarchical tables (Cheng et al., 2022; Katsis et al., 2022; Ma et al., 2024), and a complex question answering benchmark has been proposed to bridge the gap between theoretical and real-world tabular data (Wu et al., 2024), the tables in these benchmarks are not complex enough and there has been no detailed discussion of the tables.
@@ 3 RealHiTBench
In practical applications, tabular data holds significant value. Numerous previous studies have proposed a series of benchmarks  (Zhu et al., 2021; Parikh et al., 2020), which have effectively propelled research in this field. However, as application scenarios become increasingly complex and large language models (LLMs) gradually emerge as the mainstream method for table reasoning (Yang et al., 2024b), the previous benchmarks struggle to provide a comprehensive and accurate assessment of model capabilities. Hence, there is an urgent need for a new, realistic, and challenging benchmark to evaluate the progress of research methods in this field. We propose RealHiTBench focusing on complex structures and tasks. Complex tables, including hierarchical tables, are widely used in various domains (Lu et al., 2023; Katsis et al., 2022). We propose such a challenging benchmark full of complex tables and corresponding difficult tasks to assess the upper bound of the comprehension capability of LLMs, which may further inspire future research on processing tables with LLM.
@@ 3.1 Table Collection and Process
To construct a realistic and comprehensive dataset, all tables in our benchmark are raw from 13 different open platforms and cover 24 different domains like economy, society, science, and so on (shown in the Appendix A.1 and A.3). In real applications, table-related benchmarks are presented in either textual  (Cheng et al., 2022; Parikh et al., 2020) or visual  (Kim et al., 2024) manner. However, there is still no fixed format for complex table understanding. Therefore, we explore the influence of textual and visual ways. Specifically, we also explore the influence of different input formats, including LaTeX, HTML, CSV, Markdown, and PNG, whose performance is shown in Figure A.2.
@@ 3.2 Complex Structures
In actual scenarios, we often encounter tables that show structural and even semantic complexity. However, current benchmarks (Cheng et al., 2022; Lu et al., 2023) have not adequately discussed the complexity of table structure. Concentrating on complexity in tabular data, we define 5 complex structure types as following definitions. Most types of complex structure are presented in Figure 1.
@@
Speaking of the dimensions of a table, single-table scenarios cannot necessarily reflect the complexity of real-world applications, and multi-table tasks have been proposed in recent work (Wu et al., 2025). We extend the representation of multiple tables to two refined categories: (i) Explicit Multi-Table Join: We define the majority of the multi-tables that have been discussed as Explicit Multi-Table (shown in Figure 14). (ii) Implicit Multi-Table Join: However, we notice other special multi-tables in our dataset. As depicted with red box in Figure 1, there are sub-tables with the same structures, especially column headers. The implicit Multi-Table type looks not different from a normal single table, but in fact, it is in the form of multiple tables. Interestingly, this type sometimes contains additional semantics, such as comparisons, but is not easy to detect when comprehending tables.
@@
It is worth noting that the hierarchical information we consider also appears in some existing benchmarks. However, we include tables with higher complexity (see Table 8 for more details). Notably, it is necessary to clarify some differences between our work and the previous HiTab (Cheng et al., 2022) benchmark. First, HiTab pre-extracts the tree structure of hierarchical tables in JSON format, which prevents effective evaluation of whether LLMs can directly understand structural information from table inputs. Additionally, it suffers from a restricted focus on three domains, a simplistic QA task, and incomplete supervision.
@@ 3.3 Complex Tasks
In order to evaluate the model’s abilities comprehensively, we define 5 primary types following TableBench (Wu et al., 2024): Fact Checking (FC), Numerical Reasoning (NR), Data Analysis (DA), Chart Generation (CG), and Structure Comprehending (SC) (in the Table 2). More detailed subtypes (shown in Appendix B.1) are derived from the above types. Notably, we extend one type named Structure Comprehending, which is tailored for complex tables. This kind of task provides the new table by exchanging some complex parts of the source table. Then we ask LLMs the same question with two similar tabular input to evaluate the ability of LLMs to comprehend structures.
@@ 3.4 Annotation Generation
Following the previous work (Yu et al., 2023a), we select some attributes that we also agree are important as the core for question construction from what GPT-4o considers. After feeding different prompts of each question type and acquiring generated questions, we conduct manual question review according to four factors: Relevance, Completeness, Feasibility, and Clarity. After the first turn of annotation, we swap questions of each other to check the rationality.
@@
To eliminate the instability of GPT generation, we generate answers with GPT-4o based on well-designed answer prompt and innovate the pipeline: We feed a table with prompts to several models, acquiring multiple results and the frequency of occurrence of each result. Then the results are input into the G-Eval (Liu et al., 2023) module and scored for each result. Thus, final scores of each result can be figured out with initial score and frequency. The result with the highest final score becomes the candidate answer. We also conduct careful human check again that the table expert is dedicated to check answers one by one according to the question and exact table content. Furthermore, we employ a tree-based method to enhance our annotation process, finding that the automatic annotation accuracy is further improved.
@@ 3.5 Dataset Statistics
We propose RealHiTBench containing 708 tables and 3,752 questions. RealHiTBench completely focuses on discussing complex tables with such characteristics: (1) Tables in our benchmark are totally raw from available public platforms, and our benchmark totally consists of complex tables in structure or semantics. Thus, we organize 6 university students as annotators spending 150 hours each person on collecting these tables. By the way, annotators spend 480 hours each person on question answering annotations. (2) Compared to work involving complex tables such as HiTab (Cheng et al., 2022) and MultiHierTT (Zhao et al., 2022) before, the tables in our dataset are more complex (quantized in Table 8), even the median sizes of the tables are larger than tables in other works  (Wu et al., 2024). (3) We systematically define different types of complex tabular structure and classify them for further exploration. (4) We introduce a couple of challenging task types. The tasks of structure comprehending prove to be difficult for LLMs (shown in section 5.3). The above characteristics are shown in Figure 1 and Table 1.
@@ 4.1 Tree Generation
We first enable the model to self-explain the structural information of the table headers, including their meanings, scopes of influence, hierarchical structure, and relationships between headers. Referring to the previous approach (Zhao et al., 2023), we prompt the model by encoding each node of the table header as a tuples $T=(t_{1},t_{2},t_{3},t_{4})$. Specifically, the first element represents row header (R) or column header (C), along with its corresponding level. And the second and third elements represent its start and end positions, while the fourth element contains the content of the cell.
@@ 4.2 Tree-based Reasoning
Previous studies Shi et al. (2023) have shown that LLMs are often distracted by irrelevant information. To address this issue, we prompt the model to decompose question into keywords and align them with the table headers, thereby helping it accurately identify the sub-table relevant to the question.
@@ 5.1 Experimental Setup
We evaluated 25 models, including different modalities, with parameters ranging from 7B to 90B across four categories: (1) Close-source models, including GPT-o1 (OpenAI, 2024), GPT-4o (OpenAI, 2023), Deepseek-R1-API(Guo et al., 2025), Gemini-1.5-pro (Anil et al., 2023), QwQ (Team, 2024) and Doubao-1.5-pro (Team, 2025) . (2) Open-source language models, including Llama3s (Dubey et al., 2024), Qwen2.5s Yang et al. (2024a), Mistral (Jiang et al., 2023) and Deepseek-R1-Distalls (Guo et al., 2025). (3) Open-source multimodal models, including LLaVA-1.5 (Liu et al., 2024), mPLUG-owl3 (Ye et al., 2024), mPLUG-owl2 (Ye et al., 2023), and Llama3-Visions. (4) Table-oriented models, including TableGPT (Su et al., 2024), TableLLMs (Wu et al., 2024), TableLlama (Zhang et al., 2024), and Table-LLaVA (Zheng et al., 2024).
@@
For Fact Checking, Numerical Reasoning and Structure Comprehending, we refer to previous work (Zhao et al., 2022) using F1 and EM as score metric for these tasks. For chart generation, we calculate the ECR to assess code executability, extract y-axis values to compare with reference data, and compute PASS@1 to evaluate the pass rate. For Data Analysis, we calculate ROUGE-L (Lin, 2004) to evaluate the objective similarity between generated and reference answers. Inspired by G-EVAL(Liu et al., 2023), we also design an evaluation template using GPT-4o to evaluate the scores of model answers. A detailed evaluation prompt template can be found in Appendix C.3.
@@ A.3 Data Domains
We are devoted to constructing a realistic benchmark to reflect the complexity of tables in real scenarios. However, open-domain tables typically exhibit much simpler structures than the complex structures found in specific domains. Tables in specific domains commonly obtain characteristics that they contain domain-specific terms and typically not accompanied by domain-specific labeled data. Thus, it is not promising to conduct supervised fine-tuning on models for TableQA tasks Katsis et al. (2022). Therefore, we cover 24 domains (shown in Figure 6) to align with real-world application scenarios. And we maintain a balance in the number of tables under each domain as much as possible.
@@ B.1 Task types
We are devoted to evaluate the performance of LLMs in table understanding and reasoning comprehensively. And table question answering (TableQA) is one of the most common applications about tables. Thus, referring to the previous work (Wu et al., 2024), we introduce 5 primary task types and a couple of more specific subtypes. Here we explain each type to avoid ambiguity.
@@ B.2 QA Annotation
In order to display the complexity in our dataset comprehensively, we introduce LLM to assist in generating annotations. Before generating questions, we interact with GPT to acquire some inspirations about the attribute dimensions of question that are vital for constructing tasks in tabular reasoning (Yu et al., 2023b). We select some points from the GPT answers, including Complexity, Diversity, Real-World Relevance, and Length. Then, we design our prompt templates from these dimensions (shown in Table  9). Moreover, we also design different prompt templates for each task type to ensure highlighting the characteristics of each type. Eventually, we combine the prompt template with the exact question type and table content (xlsx format to simulate human intuitive vision) as input for GPT, acquiring the generated answer.
@@
Referring to the previous work, the LLM-assisted annotation is likely to lead to the bias or instability of benchmark (Felkner et al., 2024). Therefore, we set question review step to ensure the rationality of questions generated by GPT. We set four attributes of a reasonable question for annotators. Then we provide the tables and questions along with these attributes as review standard to annotators to confirm rationality or adjust the questions.

### v2

@@ 1 Introduction
In recent years, there has been a lot of work on table analysis (table_reasoning_1; TAPERA). Specifically, an increasing number of studies discuss table question answering (TableQA) tasks. (effective_distillation; table_sequence; table_recognition; Table_LLM_2; GraphOTTER). Additionally, the introduction of large language models (LLMs), including close-source models such as GPT-4o (gpt) and Gemini-1.5-Pro (gemini), and table-oriented models like TableGPT (tablegpt) further enhances models’ capability of table comprehension. Given these advancements, the rapid development of LLMs and the continuous emergence of diverse tabular data in real-world scenarios have highlighted the need for a comprehensive evaluation of LLMs’ table understanding capabilities, putting forward new requirements on the relevant benchmarks.
@@
Unfortunately, we find that the current tabular data benchmarks, such as TAT-QA (TAT-QA), TableBench (tablebench), and InfiAgent-DABench (infiAgent), largely consist of flat tables. More concretely, in such tables, each column represents an attribute, each row represents a record, and all data are stored in a simple one-dimensional format; see Figure 13 for visual demonstration. However, in practical applications, humans often organize relatively complex tables to represent multifaceted relationships between variables. Such hierarchically structured tables are popular in various domains, including economy, science, and employment, on public data platforms.
To address this issue, some benchmarks have considered hierarchical tables, such as HiTab (HiTab) and SciTab (scitab). However, these benchmarks fail to comprehensively present the LLMs’ understanding capabilities of complex table structures. For example, SciTab (scitab) and AIT-QA (AIT-QA) only consider tables of the scientific and airline domains, respectively. MM-Tab (multimodal_table_understanding) provides only image-based input, while hierarchical tables in realistic applications are also presented in a textual way. SpreadSheetBench (SpreadsheetBench) focuses on the operation of tables. HiTab (HiTab) organizes its tables with lossy JSON format, only contains fundamental QA tasks, and provides incomplete supervision. More importantly, most of these benchmarks focus primarily on relatively simple hierarchical tables, particularly those with a basic column hierarchy, where the hierarchy typically does not exceed two levels. To date, there is still a lack of dedicated benchmarks for comprehensively evaluating LLMs’ ability to understand complex tabular hierarchies.
@@ 2 Related Work
Table analysis is pivotal across numerous domains (text2analysis), while table question answering (TableQA) can effectively assess analytical capabilities (origin_tableQA). Recently, there has been a growing number of studies focusing on the ability of LLMs to comprehend tabular data (Table_LLM_1; Table_LLM_2; Table_LLM_3). However, the development of LLMs leads to the emergence of different branches. Common tabular data are in text form to be processed by LLMs (gemini; qwen2.5; llama3), while other image-based tables are suited to be processed by MLLMs (Qwen2-VL; llava). Additionally, some table-oriented LLMs for both modalities have also been proposed (tablegpt; multimodal_table_understanding). Therefore, the flourishing development of LLMs makes challenging benchmarks for tabular data increasingly important.
@@
An increasing number of benchmarks for table analysis are discussed (TabFact; fetaQA). However, a proportion of benchmarks focus more on other aspects, such as reasoning methods and supervision fine-tuning, while overlooking the structural complexity of the data, especially the table dimensions (tablebench; multimodal_table_understanding). Although some recent benchmarks have introduced the concepts of hierarchical tables (HiTab; AIT-QA; SpreadsheetBench), and a complex question answering benchmark has been proposed to bridge the gap between theoretical and real-world tabular data (tablebench), the tables in these benchmarks are not complex enough and there has been no detailed discussion of the tables.
@@ 3 RealHiTBench
In practical applications, tabular data holds significant value. Numerous previous studies have proposed a series of benchmarks  (TAT-QA; ToTTo), which have effectively propelled research in this field. However, as application scenarios become increasingly complex and large language models (LLMs) gradually emerge as the mainstream method for table reasoning (effective_distillation), the previous benchmarks struggle to provide a comprehensive and accurate assessment of model capabilities. Hence, there is an urgent need for a new, realistic, and challenging benchmark to evaluate the progress of research methods in this field. We propose RealHiTBench focusing on complex structures and tasks. Complex tables, including hierarchical tables, are widely used in various domains (scitab; AIT-QA). We propose such a challenging benchmark full of complex tables and corresponding difficult tasks to assess the upper bound of the comprehension capability of LLMs, which may further inspire future research on processing tables with LLM.
@@ 3.1 Table Collection and Process
To construct a realistic and comprehensive dataset, all tables in our benchmark are raw from 13 different open platforms and cover 24 different domains like economy, society, science, and so on (shown in the Appendix A.1 and A.3). In real applications, table-related benchmarks are presented in either textual  (HiTab; ToTTo) or visual  (TableVQA-Bench) manner. However, there is still no fixed format for complex table understanding. Therefore, we explore the influence of textual and visual ways. Specifically, we also explore the influence of different input formats, including LaTeX, HTML, CSV, Markdown, and PNG, whose performance is shown in Figure A.2.
@@ 3.2 Complex Structures
In actual scenarios, we often encounter tables that show structural and even semantic complexity. However, current benchmarks (HiTab; scitab) have not adequately discussed the complexity of table structure. Concentrating on complexity in tabular data, we define 5 complex structure types as following definitions. Most types of complex structure are presented in Figure 1.
@@
Speaking of the dimensions of a table, single-table scenarios cannot necessarily reflect the complexity of real-world applications, and multi-table tasks have been proposed in recent work (MMQA). We extend the representation of multiple tables to two refined categories: (i) Explicit Multi-Table Join: We define the majority of the multi-tables that have been discussed as Explicit Multi-Table (shown in Figure 14). (ii) Implicit Multi-Table Join: However, we notice other special multi-tables in our dataset. As depicted with red box in Figure 1, there are sub-tables with the same structures, especially column headers. The implicit Multi-Table type looks not different from a normal single table, but in fact, it is in the form of multiple tables. Interestingly, this type sometimes contains additional semantics, such as comparisons, but is not easy to detect when comprehending tables.
@@
It is worth noting that the hierarchical information we consider also appears in some existing benchmarks. However, we include tables with higher complexity (see Table 8 for more details). Notably, it is necessary to clarify some differences between our work and the previous HiTab (HiTab) benchmark. First, HiTab pre-extracts the tree structure of hierarchical tables in JSON format, which prevents effective evaluation of whether LLMs can directly understand structural information from table inputs. Additionally, it suffers from a restricted focus on three domains, a simplistic QA task, and incomplete supervision.
@@ 3.3 Complex Tasks
In order to evaluate the model’s abilities comprehensively, we define 5 primary types following TableBench (tablebench): Fact Checking (FC), Numerical Reasoning (NR), Data Analysis (DA), Chart Generation (CG), and Structure Comprehending (SC) (in the Table 2). More detailed subtypes (shown in Appendix B.1) are derived from the above types. Notably, we extend one type named Structure Comprehending, which is tailored for complex tables. This kind of task provides the new table by exchanging some complex parts of the source table. Then we ask LLMs the same question with two similar tabular input to evaluate the ability of LLMs to comprehend structures.
@@ 3.4 Annotation Generation
Following the previous work (Attributes), we select some attributes that we also agree are important as the core for question construction from what GPT-4o considers. After feeding different prompts of each question type and acquiring generated questions, we conduct manual question review according to four factors: Relevance, Completeness, Feasibility, and Clarity. After the first turn of annotation, we swap questions of each other to check the rationality.
@@
To eliminate the instability of GPT generation, we generate answers with GPT-4o based on well-designed answer prompt and innovate the pipeline: We feed a table with prompts to several models, acquiring multiple results and the frequency of occurrence of each result. Then the results are input into the G-Eval (G-EVAL) module and scored for each result. Thus, final scores of each result can be figured out with initial score and frequency. The result with the highest final score becomes the candidate answer. We also conduct careful human check again that the table expert is dedicated to check answers one by one according to the question and exact table content. Furthermore, we employ a tree-based method to enhance our annotation process, finding that the automatic annotation accuracy is further improved.
@@ 3.5 Dataset Statistics
We propose RealHiTBench containing 708 tables and 3,752 questions. RealHiTBench completely focuses on discussing complex tables with such characteristics: (1) Tables in our benchmark are totally raw from available public platforms, and our benchmark totally consists of complex tables in structure or semantics. Thus, we organize 6 university students as annotators spending 150 hours each person on collecting these tables. By the way, annotators spend 480 hours each person on question answering annotations. (2) Compared to work involving complex tables such as HiTab (HiTab) and MultiHierTT (MULTIHIERTT) before, the tables in our dataset are more complex (quantized in Table 8), even the median sizes of the tables are larger than tables in other works  (tablebench). (3) We systematically define different types of complex tabular structure and classify them for further exploration. (4) We introduce a couple of challenging task types. The tasks of structure comprehending prove to be difficult for LLMs (shown in section 5.3). The above characteristics are shown in Figure 1 and Table 1.
@@ 4.1 Tree Generation
We first enable the model to self-explain the structural information of the table headers, including their meanings, scopes of influence, hierarchical structure, and relationships between headers. Referring to the previous approach (TableTree), we prompt the model by encoding each node of the table header as a tuples $T=(t_{1},t_{2},t_{3},t_{4})$. Specifically, the first element represents row header (R) or column header (C), along with its corresponding level. And the second and third elements represent its start and end positions, while the fourth element contains the content of the cell.
@@ 4.2 Tree-based Reasoning
Previous studies Irrelevant have shown that LLMs are often distracted by irrelevant information. To address this issue, we prompt the model to decompose question into keywords and align them with the table headers, thereby helping it accurately identify the sub-table relevant to the question.
@@ 5.1 Experimental Setup
We evaluated 25 models, including different modalities, with parameters ranging from 7B to 90B across four categories: (1) Close-source models, including GPT-o1 (GPT-o1), GPT-4o (gpt), Deepseek-R1-API(deepseek), Gemini-1.5-pro (gemini), QwQ (QWQ) and Doubao-1.5-pro (doubao) . (2) Open-source language models, including Llama3s (llama3), Qwen2.5s qwen2.5, Mistral (mistral) and Deepseek-R1-Distalls (deepseek). (3) Open-source multimodal models, including LLaVA-1.5 (llava), mPLUG-owl3 (mPLUG-Owl3), mPLUG-owl2 (mPLUG-Owl2), and Llama3-Visions. (4) Table-oriented models, including TableGPT (tablegpt), TableLLMs (tablebench), TableLlama (tablellama), and Table-LLaVA (multimodal_table_understanding).
@@
For Fact Checking, Numerical Reasoning and Structure Comprehending, we refer to previous work (MULTIHIERTT) using F1 and EM as score metric for these tasks. For chart generation, we calculate the ECR to assess code executability, extract y-axis values to compare with reference data, and compute PASS@1 to evaluate the pass rate. For Data Analysis, we calculate ROUGE-L (rouge) to evaluate the objective similarity between generated and reference answers. Inspired by G-EVAL(G-EVAL), we also design an evaluation template using GPT-4o to evaluate the scores of model answers. A detailed evaluation prompt template can be found in Appendix C.3.
@@ A.3 Data Domains
We are devoted to constructing a realistic benchmark to reflect the complexity of tables in real scenarios. However, open-domain tables typically exhibit much simpler structures than the complex structures found in specific domains. Tables in specific domains commonly obtain characteristics that they contain domain-specific terms and typically not accompanied by domain-specific labeled data. Thus, it is not promising to conduct supervised fine-tuning on models for TableQA tasks AIT-QA. Therefore, we cover 24 domains (shown in Figure 6) to align with real-world application scenarios. And we maintain a balance in the number of tables under each domain as much as possible.
@@ B.1 Task types
We are devoted to evaluate the performance of LLMs in table understanding and reasoning comprehensively. And table question answering (TableQA) is one of the most common applications about tables. Thus, referring to the previous work (tablebench), we introduce 5 primary task types and a couple of more specific subtypes. Here we explain each type to avoid ambiguity.
@@ B.2 QA Annotation
In order to display the complexity in our dataset comprehensively, we introduce LLM to assist in generating annotations. Before generating questions, we interact with GPT to acquire some inspirations about the attribute dimensions of question that are vital for constructing tasks in tabular reasoning (attribute). We select some points from the GPT answers, including Complexity, Diversity, Real-World Relevance, and Length. Then, we design our prompt templates from these dimensions (shown in Table  9). Moreover, we also design different prompt templates for each task type to ensure highlighting the characteristics of each type. Eventually, we combine the prompt template with the exact question type and table content (xlsx format to simulate human intuitive vision) as input for GPT, acquiring the generated answer.
@@
Referring to the previous work, the LLM-assisted annotation is likely to lead to the bias or instability of benchmark (LLM_is_not_annotator). Therefore, we set question review step to ensure the rationality of questions generated by GPT. We set four attributes of a reasonable question for annotators. Then we provide the tables and questions along with these attributes as review standard to annotators to confirm rationality or adjust the questions.